# Nonreciprocal entanglement in cavity magnomechanics exploiting chiral cavity-magnon coupling

### v3

@@ Abstract
We show how to achieve nonreciprocal quantum entanglement in a cavity magnomechanical system by exploiting the chiral cavity-magnon coupling. The system consists of a magnon mode, a mechanical vibration mode, and two degenerate counter-propagating microwave cavity modes in a torus-shaped cavity. We show that nonreciprocal stationary microwave-magnon and -phonon bipartite entanglements and photon-magnon-phonon tripartite entanglement can be achieved by respectively driving different circulating cavity modes that hold a chiral coupling to the magnon mode. The nonreciprocal entanglements are shown to be robust against various experimental imperfections. We specifically show how such nonreciprocal entanglement can realize the channel multiplexing quantum teleportation. The work may find promising applications of the cavity magnomechanical systems in noise-tolerant quantum processing, channel multiplexing quantum teleportation, and chiral magnonic quantum networks.
I Introduction
@@
Hybrid systems based on collective spin excitations (magnons) in magnetic materials, e.g., yttrium iron garnet (YIG), have attracted considerable attention in the past decade. This is due to many excellent properties of the magnonic system, such as a large frequency tunability, a low dissipation rate, and an excellent ability to coherently couple with a variety of physical systems, including microwave photons Huebl2013 ; Tabuchi2014 ; Zhang2014 , optical photons Osada2016 ; Osada2018 ; ZhangX2016 ; Haigh2016 , vibration phonons Zhang2016 ; Li2018 ; Potts2021 ; Shen2022 , superconducting qubits Lachance-Quirion2017 ; Lachance-Quirion2020 ; Tabuchi2015 ; Xu2023 , nitrogen-vacancy spins Hei2023 ; WangY2023 , etc. As a subfield of hybrid magnonics, the couplings among microwave cavity photons, magnons, and magnetostriction-induced vibration phonons form the system of cavity magnomechanics (CMM) Zuo2023 . The CMM system provides a hybrid platform not only for studying strong interactions between light and matter, e.g., the triple magnon-photon-phonon strong coupling has been achieved Shen23 ; for studying rich nonlinear effects, such as Kerr-type nonlinearities Shen2022 , magnonic frequency combs Xiong23FR ; Dong23 , synchronization Cheng23 , non-Gaussianity and self-sustaining dynamics Wen23 , but also for preparing many different kinds of quantum states, including entangled states Li2018 ; Li2019A ; Li2020 ; Li2021B , squeezed states Li2019B ; ZhangW2021 ; Qian2023B ; Li2023 , quantum steering Tan2019 ; ZhangW2022 ; Chen2021 ; Tan2023 , and quantum ground states of mechanical motion Ding20 ; Lu21 ; Asjad23 . These quantum states find potential applications in macroscopic quantum phenomena, quantum information processing, and quantum sensing Zuo2023 ; Lachance-Quirion2019 ; Yuan2022 .
@@
Among many burgeoning research fields in advanced photonics, the chiral quantum optics Lodahl2017 , which exploits the new approaches and physical systems with chiral light-matter interaction Petersen2014 ; Kim2015 ; Dong2015 ; Gong2018 , finds many promising applications ranging from light technology to quantum information science Sollner2015 ; Mahmoodian2016 . The chiral interaction can be achieved, e.g., by coupling the spin-momentum-locked light to quantum systems with polarization-dependent dipole transitions, such as atoms Mitsch2014 ; Pucher2022 and quantum dots Coles2016 . It is of significance to extend the advances in chiral optics to the microwave domain. In particular, along with the fast development of cavity magnonics Lachance-Quirion2019 ; Yuan2022 ; Bauer22 , the chiral microwave-magnon coupling has recently been proposed Yu2020 and demonstrated Bourhill2023 by employing a torus-shaped microwave cavity. Besides, it has also been realized in the substrate-integrated-waveguide cavity Zhang2020 .
@@
Here, we propose to achieve nonreciprocal quantum entanglement in CMM by exploiting the chiral cavity-magnon coupling. We note that although nonreciprocal classical phenomena have been widely studied in various systems, nonreciprocal quantum effects have been much less explored. After the pioneering works of this kind Huang2018 ; Jiao2020 , many nonreciprocal quantum protocols have been offered in cavity magnonics and the related field of optomechanics, including nonreciprocal entanglement Jiao2020 ; Yang2020 ; HJing22 ; Ren2022 ; Yang2023 ; Chakraborty2023 ; Chen2023B ; Zheng2023 ; LYe ; Liu2023 ; Zheng2024 , quantum steering Tan2019 ; ZhangW2022 ; Chen2021 ; Tan2023 ; Yang2021 ; Tan2021 ; Zhan2022 ; Guan2022 ; Zhong2023 , squeezing Xie2023 ; Guo2023 , photon Huang2018 ; LiB2019 ; Xu2020 ; XieH2022 and magnon You22 blockade, phonon Jiang2018 ; Xu2021 and magnon XuY2021 ; Huang2022 lasing, single-photon scattering Hafezi2012 ; Xu2017 ; Ren2022A ; Xie2023A ; Xiong23 , etc. Our proposal is based on the CMM system combining with the chiral cavity-magnon coupling realized in a torus-shaped microwave cavity. Specifically, we consider two degenerate counter-propagating microwave cavity modes, a magnon (Kittel) mode, and a mechanical vibration mode. The chiral cavity-magnon coupling is achieved by placing the YIG sphere on a special line in the microwave cavity, where the Kittel mode couples solely to a specific circulating cavity mode and decouples from the other counter-propagating mode, such that the time-reversal symmetry between the two degenerate modes is broken Yu2020 ; Bourhill2023 . For the case of the cavity-magnon coupling being established, the effective magnomechanical coupling is greatly enhanced when the associated circulating cavity mode is strongly driven. When the drive field is red-detuned from the magnon mode and the magnomechanical coupling becomes sufficiently strong, the magnon-phonon entanglement can be created, which further distributes to the circulating microwave mode, yielding the microwave-magnon and -phonon entanglements and the photon-magnon-phonon tripartite entanglement. For the case where the cavity-magnon coupling is absent, the magnomechanical subsystem is essentially decoupled from the (strongly driven) cavity. In this case, the magnomechanical dispersive coupling is too weak to generate any entanglement in the system. Consequently, the chiral cavity-magnon coupling leads to the nonreciprocity of the microwave-magnon and -phonon entanglements and the tripartite entanglement when the two counter-propagating microwave cavity modes are respectively driven, e.g., via a waveguide. We further analyze the effects of various imperfections in a real experiment, e.g., the backscattering induced coupling between the two degenerate modes and the inaccurate position of the YIG sphere in the cavity, which causes simultaneous couplings with the two degenerate modes, and find that the nonreciprocity of the entanglements is robust against these imperfections.
@@
The paper is organized as follows. In Sec. II, we introduce the system, provide its Hamiltonian and the corresponding quantum Langevin equations (QLEs), and show how to obtain steady-state entanglements of the system. We then present the results of the nonreciprocal entanglements achieved in the ideal situation in Sec. III, and analyze the effects of some major experimental imperfections in Sec. IV. We further discuss many potential applications of the proposed nonreciprocal entanglements in Sec. V, and draw the conclusions in Sec. VI.
@@
II The Model
The CMM system under study consists of a torus-shaped metallic microwave cavity Yu2020 ; Bourhill2023 and a ferrimagnetic sample, e.g., a YIG sphere, that is placed in the cavity; see Fig. 1(a). The microwave cavity holds a pair of degenerate counter-propagating transverse electric (TE) microwave modes, which couple to a microwave waveguide through which a specific circulating mode is driven. The YIG sphere supports both a magnon (spin wave) mode and a mechanical vibration mode induced by the magnetostrictive interaction, which is a dispersive interaction for a large-size YIG sphere and provides necessary nonlinearity for creating entanglement in our protocol Zuo2023 . The magnon mode further couples to the microwave cavity mode via the magnetic-dipole interaction. Specifically, we consider the magnon mode to be the Kittel mode, which corresponds to the uniform precession of all spins in the YIG sphere. The Hamiltonian of the system reads
@@ II The Model
where $a_{\circlearrowright}$, $a_{\circlearrowleft}$ and $m$ ($a_{\circlearrowright}^{\dagger}$, $a_{\circlearrowleft}^{\dagger}$ and $m^{\dagger}$) denote the annihilation (creation) operators of the clockwise (CW) and counter-clockwise (CCW) circulating microwave cavity modes and the Kittel mode, respectively, and $q$ and $p$ are the dimensionless position and momentum of the mechanical motion, satisfying $[q,p]=i$. $\omega_{a}$, $\omega_{m}$, and $\omega_{b}$ are the resonance frequencies of the two degenerate cavity modes, the magnon mode, and the mechanical mode, respectively. The cavity-magnon coupling $g_{\circlearrowright}$ ($g_{\circlearrowleft}$) describes the interaction strength between the magnons and the CW (CCW) circulating microwave photons. The single-magnon magnomechanical coupling rate $g_{m}$ is typically weak, but the effective magnomechanical coupling $G_{m}$ can be greatly enhanced by directly driving the magnon mode Shen23 , or driving the microwave cavity via the cavity-magnon excitation-exchange interaction Zhang2016 ; Potts2021 ; Shen2022 . The last term $H_{d}/\hbar=iE(a_{l}^{\dagger}e^{-i\omega_{0}t}-a_{l}e^{i\omega_{0}t})$ describes the microwave driving of the CW ($l=\circlearrowright$) or CCW ($l=\circlearrowleft$) circulating cavity mode, where $E=\sqrt{2\kappa_{a,e}P_{0}/\hbar\omega_{0}}$ is the coupling strength between the cavity mode and the drive field, with $\kappa_{a,e}$ being the external dissipation rate of the cavity and $P_{0}$ ($\omega_{0}$) the power (frequency) of the drive field.
The symmetry between the CW and CCW circulating microwave modes in coupling with the Kittel mode can be manipulated to be broken by placing the YIG sphere on a special line in the torus-shaped cavity, which leads to the chiral cavity-magnon coupling. This is due to the fact that the Kittel mode magnetization precesses anticlockwise around the effective magnetic field and couples to the microwave photons with the same polarization, and that the polarization of some cavity modes depends on their propagation direction at special lines in the cavity Yu2020 . To be specific, the Kittel mode only couples to the right-handed circularly polarized magnetic field with respect to the $z$ axis. For a TE mode, this corresponds to a CW (CCW) mode for the YIG sphere being positioned on the special line of, e.g., $r=r_{+}$ ($r=r_{-}$), cf. Fig. 1(a), where $r$ is the radial coordinate Bourhill2023 . Accordingly, the Kittel mode decouples to the counter-propagating cavity mode on the same circular line. As a result, the chiral cavity-magnon coupling is achieved. The chiral coupling shows good tunability: one can realize the cavity-magnon couplings taking either $g_{\circlearrowright}>0$ and $g_{\circlearrowleft}=0$, or $g_{\circlearrowleft}>0$ and $g_{\circlearrowright}=0$, by placing the YIG sphere on a specific orbit.
@@
where $\Delta_{\circlearrowright(\circlearrowleft)}\equiv\Delta_{a}=\omega_{a}-\omega_{0}$, $\Delta_{m}=\omega_{m}-\omega_{0}$, $\kappa_{m}$ ($\gamma_{b}$) is the dissipation rate of the magnon (mechanical) mode, and $\kappa_{a}=\kappa_{a,i}+\kappa_{a,e}$ represents the total dissipation rate of the microwave cavity, with $\kappa_{a,i}$ being the internal dissipation rate. Here, $k^{\mathrm{in}}$ ($k=a_{\circlearrowright}$, $a_{\circlearrowleft}$, $m$) denote the zero-mean input noises of the CW, CCW, and magnon modes, respectively. They obey the following correlation functions: $\langle k^{\mathrm{in}}(t)k^{\mathrm{in}\dagger}(t^{\prime})\rangle=[N_{k}(\omega_{k})+1]\delta(t-t^{\prime})$ and $\langle k^{\mathrm{in}\dagger}(t)k^{\mathrm{in}}(t^{\prime})\rangle=N_{k}(\omega_{k})\delta(t-t^{\prime})$. The Brownian noise operator $\xi(t)$ is for the mechanical oscillator, which is intrinsically non-Markovian. Nevertheless, for a high mechanical quality factor $Q_{b}=\omega_{b}/\gamma_{b}\gg 1$, the Markovian approximation can be made Giova2001 , leading $\xi(t)$ to be $\delta$-correlated: $\langle\xi(t)\xi(t^{\prime})+\xi(t^{\prime})\xi(t)\rangle/2\simeq\gamma_{b}[2N_{b}(\omega_{b})+1]\delta(t-t^{\prime})$. In the above noise correlations, $N_{j}(\omega_{j})=[\mathrm{exp}(\hbar\omega_{j}/k_{B}T)-1]^{-1}$ ($j=a$, $m$, $b$) is the mean thermal excitation number of the corresponding mode at bath temperature $T$, with $k_{B}$ being the Boltzmann constant.
The generation of entanglement in the present system requires a sufficiently strong magnomechanical coupling strength Li2018 . To this end, we strongly drive the CW cavity mode via a microwave waveguide (Fig. 1(a)). Due to the cavity-magnon excitation-exchange interaction, the magnon mode also gets effectively pumped, leading to $|\langle m\rangle|\gg 1$. This allows us to linearize the system dynamics around the large average values by writing each operator as $O=\langle O\rangle+\delta O$ ($O\,{=}\,a_{\circlearrowright},a_{\circlearrowleft},m,q,p$) and neglecting small second-order fluctuation terms. Consequently, we obtain a set of linearized QLEs for the quantum fluctuations of the system $(\delta a_{\circlearrowright},\delta a_{\circlearrowleft},\delta m,\delta q,\delta p)$, which can be rewritten in the quadrature form as follows
@@
$A{=}\begin{pmatrix}-\kappa_{a}&\Delta_{\circlearrowright}&0&0&0&g_{\circlearrowright}&0&0\\ -\Delta_{\circlearrowright}&-\kappa_{a}&0&0&-g_{\circlearrowright}&0&0&0\\ 0&0&-\kappa_{a}&\Delta_{\circlearrowleft}&0&0&0&0\\ 0&0&-\Delta_{\circlearrowleft}&-\kappa_{a}&0&0&0&0\\ 0&g_{\circlearrowright}&0&0&-\kappa_{m}&\tilde{\Delta}_{m}&\mathrm{Im}\,G_{m}&0\\ -g_{\circlearrowright}&0&0&0&-\tilde{\Delta}_{m}&-\kappa_{m}&{-}\mathrm{Re}\,G_{m}&0\\ 0&0&0&0&0&0&0&\omega_{b}\\ 0&0&0&0&{-}\mathrm{Re}\,G_{m}&{-}\mathrm{Im}\,G_{m}&-\omega_{b}&-\gamma_{b}\end{pmatrix},$ (4)
@@
$\langle m\rangle=\frac{-ig_{\circlearrowright}E}{g_{\circlearrowright}^{2}+(\kappa_{a}+i\Delta_{a})(\kappa_{m}+i\tilde{\Delta}_{m})}.$ (5)
@@
Because of the linearized dynamics and the Gaussian nature of the input noises, the steady state of the quantum fluctuations of the system retains a Gaussian state, which can be described by an $8\times 8$ covariance matrix (CM) $V$, with its entries defined as $V_{ij}=\langle u_{i}(t)u_{j}(t^{\prime})+u_{j}(t^{\prime})u_{i}(t)\rangle/2$ ($i,j\in\{1,2,...,8\}$). At the steady state, the CM can be conveniently achieved by solving the Lyapunov equation Vitali2007
@@
where $D=\mathrm{diag}[\kappa_{a}(2N_{a}+1),\kappa_{a}(2N_{a}+1),\kappa_{a}(2N_{a}+1),\kappa_{a}(2N_{a}+1),$ $\kappa_{m}(2N_{m}+1),\kappa_{m}(2N_{m}+1),0,\gamma_{b}(2N_{b}+1)]$ is the diffusion matrix, which is defined via $D_{ij}\delta(t-t^{\prime})=\langle n_{i}(t)n_{j}(t^{\prime})+n_{j}(t^{\prime})n_{i}(t)\rangle/2$. When the CM is obtained, we adopt the logarithmic negativity $E_{N}$ to calculate the degree of any bipartite entanglement of the system, which is defined as Adesso2004
@@
III nonreciprocal entanglements in the ideal case

### v4

# 1]organization=School of Physics, Zhejiang Key Laboratory of Micro-Nano Quantum Chips and Quantum Control, and State Key Laboratory for Extreme Photonics and Instrumentation, city=Hangzhou, postcode=310027, state=Zhejiang, country=China

@@ Abstract
We propose a new mechanism to achieve nonreciprocal quantum entanglement in a cavity magnomechanical system by exploiting the chiral cavity-magnon coupling. The system consists of a magnon mode, a mechanical vibration mode, and two degenerate counter-propagating microwave cavity modes in a torus-shaped cavity. We show that nonreciprocal stationary microwave-magnon and -phonon bipartite entanglements and photon-magnon-phonon tripartite entanglement can be achieved by respectively driving different circulating cavity modes that hold a chiral coupling to the magnon mode. The nonreciprocal entanglements are shown to be robust against various experimental imperfections. We specifically show how such nonreciprocal entanglement can realize the channel multiplexing quantum teleportation from a microwave field to a solid-state magnon mode. The work may find promising applications of the cavity magnomechanical systems in noise-tolerant quantum processing, channel multiplexing quantum teleportation, and chiral magnonic quantum networks.
keywords:
[…]
1 Introduction
@@
Hybrid systems based on collective spin excitations (magnons) in magnetic materials, e.g., yttrium iron garnet (YIG), have attracted considerable attention in the past decade. This is due to many excellent properties of the magnonic system, such as a large frequency tunability, a low dissipation rate, and an excellent ability to coherently couple with a variety of physical systems, including microwave photons [1, 2, 3], optical photons [4, 7, 5, 6], vibration phonons [8, 9, 10, 11], superconducting qubits [13, 14, 12, 15], nitrogen-vacancy spins [16, 17], etc. As a subfield of hybrid magnonics, the couplings among microwave cavity photons, magnons, and magnetostriction-induced vibration phonons form the system of cavity magnomechanics (CMM) [18]. The CMM system provides a hybrid platform not only for studying strong interactions between light and matter, e.g., the triple magnon-photon-phonon strong coupling has been achieved [19]; for studying rich nonlinear effects, such as Kerr-type nonlinearities [11], magnonic frequency combs [20, 21], synchronization [22], non-Gaussianity and self-sustaining dynamics [23], but also for preparing many different kinds of quantum states, including entangled states [9, 24, 25, 26], squeezed states [27, 28, 30, 29], quantum steering [31, 33, 32, 34], and quantum ground states of mechanical motion [35, 36, 37]. These quantum states find potential applications in macroscopic quantum phenomena, quantum information processing, and quantum sensing [18, 38, 39].
@@
Among many burgeoning research fields in advanced photonics, the chiral quantum optics [40], which exploits the new approaches and physical systems with chiral light-matter interaction [41, 42, 43, 44], finds many promising applications ranging from light technology to quantum information science [45, 46]. The chiral interaction can be achieved, e.g., by coupling the spin-momentum-locked light to quantum systems with polarization-dependent dipole transitions, such as atoms [47, 48] and quantum dots [49]. It is of significance to extend the advances in chiral optics to the microwave domain. In particular, along with the fast development of cavity magnonics [38, 39, 50], the chiral microwave-magnon coupling has recently been proposed [51] and demonstrated [52] by employing a torus-shaped microwave cavity. Besides, it has also been realized in the substrate-integrated-waveguide cavity [53].
@@
Here, we propose to achieve nonreciprocal quantum entanglement in CMM by exploiting the chiral cavity-magnon coupling. We note that although nonreciprocal classical phenomena have been widely studied in various systems, nonreciprocal quantum effects have been much less explored. After the pioneering works of this kind [54, 55], many nonreciprocal quantum protocols have been offered in cavity magnonics and the related field of optomechanics, including nonreciprocal entanglement [55, 56, 57, 59, 62, 61, 60, 63, 65, 58, 64], quantum steering [31, 33, 32, 34, 66, 67, 68, 69, 70], squeezing [71, 72], photon [54, 73, 74, 75] and magnon [76] blockade, phonon [77, 78] and magnon [79, 80] lasing, single-photon scattering [81, 82, 83, 84, 85], etc. We note that most nonreciprocal entanglement proposals in cavity magnonics are based on a spinning microwave resonator configuration. However, unlike optical microcavities, it is still experimentally challenging to realize high-speed rotation of a large-size microwave cavity. Here, our proposal is based on the CMM system combining with the chiral cavity-magnon coupling realized in a torus-shaped microwave cavity, which has been demonstrated in the related experimental study [52]. Specifically, we consider two degenerate counter-propagating microwave cavity modes, a magnon (Kittel) mode, and a mechanical vibration mode. The chiral cavity-magnon coupling is achieved by placing the YIG sphere on a special line in the microwave cavity, where the Kittel mode couples solely to a specific circulating cavity mode and decouples from the other counter-propagating mode, such that the time-reversal symmetry between the two degenerate modes is broken [51, 52]. For the case of the cavity-magnon coupling being established, the effective magnomechanical coupling is greatly enhanced when the associated circulating cavity mode is strongly driven. When the drive field is red-detuned from the magnon mode and the magnomechanical coupling becomes sufficiently strong, the magnon-phonon entanglement can be created, which further distributes to the circulating microwave mode, yielding the microwave-magnon and -phonon entanglements and the photon-magnon-phonon tripartite entanglement. For the case where the cavity-magnon coupling is absent, the magnomechanical subsystem is essentially decoupled from the (strongly driven) cavity. In this case, the magnomechanical dispersive coupling is too weak to generate any entanglement in the system. Consequently, the chiral cavity-magnon coupling leads to the nonreciprocity of the microwave-magnon and -phonon entanglements and the tripartite entanglement when the two counter-propagating microwave cavity modes are respectively driven, e.g., via a waveguide. We further analyze the effects of various imperfections in a real experiment, e.g., the backscattering induced coupling between the two degenerate modes and the inaccurate position of the YIG sphere in the cavity, which causes simultaneous couplings with the two degenerate modes, and find that the nonreciprocity of the entanglements is robust against these imperfections. We specifically show how a channel multiplexing quantum teleportation is possible based on the generated nonreciprocal entanglement, which manifests the potential application of our work in quantum information processing and chiral quantum networks.
@@
The paper is organized as follows. In Sec. 2, we introduce the system, provide its Hamiltonian and the corresponding quantum Langevin equations (QLEs), and show how to obtain steady-state entanglements of the system. We then present the results of the nonreciprocal entanglements achieved in the ideal situation in Sec. 3, and analyze the effects of some major experimental imperfections in Sec. 4. We further discuss many potential applications of the proposed nonreciprocal entanglements in Sec. 5, and draw the conclusions in Sec. 6.
@@
2 The Model
The CMM system under study consists of a torus-shaped metallic microwave cavity [51, 52] and a ferrimagnetic sample, e.g., a YIG sphere, that is placed in the cavity; see Fig. 1(a). The microwave cavity holds a pair of degenerate counter-propagating transverse electric (TE) microwave modes, which couple to a microwave waveguide through which a specific circulating mode is driven. The YIG sphere supports both a magnon (spin wave) mode and a mechanical vibration mode induced by the magnetostrictive interaction, which is a dispersive interaction for a large-size YIG sphere and provides necessary nonlinearity for creating entanglement in our protocol [18]. The magnon mode further couples to the microwave cavity mode via the magnetic-dipole interaction. Specifically, we consider the magnon mode to be the Kittel mode, which corresponds to the uniform precession of all spins in the YIG sphere. The Hamiltonian of the system reads
@@ 2 The Model
where $a_{\circlearrowright}$, $a_{\circlearrowleft}$ and $m$ ($a_{\circlearrowright}^{\dagger}$, $a_{\circlearrowleft}^{\dagger}$ and $m^{\dagger}$) denote the annihilation (creation) operators of the clockwise (CW) and counter-clockwise (CCW) circulating microwave cavity modes and the Kittel mode, respectively, and $q$ and $p$ are the dimensionless position and momentum of the mechanical motion, satisfying $[q,p]=i$. $\omega_{a}$, $\omega_{m}$, and $\omega_{b}$ are the resonance frequencies of the two degenerate cavity modes, the magnon mode, and the mechanical mode, respectively. The cavity-magnon coupling $g_{\circlearrowright}$ ($g_{\circlearrowleft}$) describes the interaction strength between the magnons and the CW (CCW) circulating microwave photons. The single-magnon magnomechanical coupling rate $g_{m}$ is typically weak, but the effective magnomechanical coupling $G_{m}$ can be greatly enhanced by directly driving the magnon mode [19], or driving the microwave cavity via the cavity-magnon excitation-exchange interaction [8, 10, 11]. The last term $H_{d}/\hbar=iE(a_{l}^{\dagger}e^{-i\omega_{0}t}-a_{l}e^{i\omega_{0}t})$ describes the microwave driving of the CW ($l=\circlearrowright$) or CCW ($l=\circlearrowleft$) circulating cavity mode, where $E=\sqrt{2\kappa_{a,e}P_{0}/\hbar\omega_{0}}$ is the coupling strength between the cavity mode and the drive field, with $\kappa_{a,e}$ being the external dissipation rate of the cavity and $P_{0}$ ($\omega_{0}$) the power (frequency) of the drive field.
The symmetry between the CW and CCW circulating microwave modes in coupling with the Kittel mode can be manipulated to be broken by placing the YIG sphere on a special line in the torus-shaped cavity, which leads to the chiral cavity-magnon coupling. This is due to the fact that the Kittel mode magnetization precesses anticlockwise around the effective magnetic field and couples to the microwave photons with the same polarization, and that the polarization of some cavity modes depends on their propagation direction at special lines in the cavity [51]. To be specific, the Kittel mode only couples to the right-handed circularly polarized magnetic field with respect to the $z$ axis. For a TE mode, this corresponds to a CW (CCW) mode for the YIG sphere being positioned on the special line of, e.g., $r=r_{+}$ ($r=r_{-}$), cf. Fig. 1(a), where $r$ is the radial coordinate [52]. Accordingly, the Kittel mode decouples to the counter-propagating cavity mode on the same circular line. As a result, the chiral cavity-magnon coupling is achieved. The chiral coupling shows good tunability: one can realize the cavity-magnon couplings taking either $g_{\circlearrowright}>0$ and $g_{\circlearrowleft}=0$, or $g_{\circlearrowleft}>0$ and $g_{\circlearrowright}=0$, by placing the YIG sphere on a specific orbit.
@@
where $\Delta_{\circlearrowright(\circlearrowleft)}\equiv\Delta_{a}=\omega_{a}-\omega_{0}$, $\Delta_{m}=\omega_{m}-\omega_{0}$, $\kappa_{m}$ ($\gamma_{b}$) is the dissipation rate of the magnon (mechanical) mode, and $\kappa_{a}=\kappa_{a,i}+\kappa_{a,e}$ represents the total dissipation rate of the microwave cavity, with $\kappa_{a,i}$ being the internal dissipation rate. Here, $k^{\mathrm{in}}$ ($k=a_{\circlearrowright}$, $a_{\circlearrowleft}$, $m$) denote the zero-mean input noises of the CW, CCW, and magnon modes, respectively. They obey the following correlation functions: $\langle k^{\mathrm{in}}(t)k^{\mathrm{in}\dagger}(t^{\prime})\rangle=[N_{k}(\omega_{k})+1]\delta(t-t^{\prime})$ and $\langle k^{\mathrm{in}\dagger}(t)k^{\mathrm{in}}(t^{\prime})\rangle=N_{k}(\omega_{k})\delta(t-t^{\prime})$. The Brownian noise operator $\xi(t)$ is for the mechanical oscillator, which is intrinsically non-Markovian. Nevertheless, for a high mechanical quality factor $Q_{b}=\omega_{b}/\gamma_{b}\gg 1$, the Markovian approximation can be made [86], leading $\xi(t)$ to be $\delta$-correlated: $\langle\xi(t)\xi(t^{\prime})+\xi(t^{\prime})\xi(t)\rangle/2\simeq\gamma_{b}[2N_{b}(\omega_{b})+1]\delta(t-t^{\prime})$. In the above noise correlations, $N_{j}(\omega_{j})=[\mathrm{exp}(\hbar\omega_{j}/k_{B}T)-1]^{-1}$ ($j=a$, $m$, $b$) is the mean thermal excitation number of the corresponding mode at bath temperature $T$, with $k_{B}$ being the Boltzmann constant.
The generation of entanglement in the present system requires a sufficiently strong magnomechanical coupling strength. To this end, we strongly drive the CW cavity mode via a microwave waveguide (Fig. 1(a)). Due to the cavity-magnon excitation-exchange interaction, the magnon mode also gets effectively pumped, leading to $|\langle m\rangle|\gg 1$. This allows us to linearize the system dynamics around the large average values by writing each operator as $O=\langle O\rangle+\delta O$ ($O\,{=}\,a_{\circlearrowright},a_{\circlearrowleft},m,q,p$) and neglecting small second-order fluctuation terms. Consequently, we obtain a set of linearized QLEs for the quantum fluctuations of the system $(\delta a_{\circlearrowright},\delta a_{\circlearrowleft},\delta m,\delta q,\delta p)$, which can be rewritten in the quadrature form as follows
@@
$\begin{split}&A{=}\\ &\begin{pmatrix}-\kappa_{a}&\Delta_{\circlearrowright}&0&0&0&g_{\circlearrowright}&0&0\\ -\Delta_{\circlearrowright}&-\kappa_{a}&0&0&-g_{\circlearrowright}&0&0&0\\ 0&0&-\kappa_{a}&\Delta_{\circlearrowleft}&0&0&0&0\\ 0&0&-\Delta_{\circlearrowleft}&-\kappa_{a}&0&0&0&0\\ 0&g_{\circlearrowright}&0&0&-\kappa_{m}&\tilde{\Delta}_{m}&\mathrm{Im}G_{m}&0\\ -g_{\circlearrowright}&0&0&0&-\tilde{\Delta}_{m}&-\kappa_{m}&-\mathrm{Re}G_{m}&0\\ 0&0&0&0&0&0&0&\omega_{b}\\ 0&0&0&0&-\mathrm{Re}G_{m}&-\mathrm{Im}G_{m}&-\omega_{b}&-\gamma_{b}\end{pmatrix},\end{split}$ (4)
@@
$\langle m\rangle=\frac{-ig_{\circlearrowright}}{\kappa_{m}+i\tilde{\Delta}_{m}}\langle a_{\circlearrowright}\rangle=\frac{-ig_{\circlearrowright}E}{g_{\circlearrowright}^{2}+(\kappa_{a}+i\Delta_{a})(\kappa_{m}+i\tilde{\Delta}_{m})},$ (5)
@@
Because of the linearized dynamics and the Gaussian nature of the input noises, the steady state of the quantum fluctuations of the system retains a Gaussian state, which can be described by an $8\times 8$ covariance matrix (CM) $V$, with its entries defined as $V_{ij}=\langle u_{i}(t)u_{j}(t^{\prime})+u_{j}(t^{\prime})u_{i}(t)\rangle/2$ ($i,j\in\{1,2,...,8\}$). At the steady state, the CM can be conveniently achieved by solving the Lyapunov equation [87]
@@
where $D=\mathrm{diag}[\kappa_{a}(2N_{a}+1),\kappa_{a}(2N_{a}+1),\kappa_{a}(2N_{a}+1),\kappa_{a}(2N_{a}+1),$ $\kappa_{m}(2N_{m}+1),\kappa_{m}(2N_{m}+1),0,\gamma_{b}(2N_{b}+1)]$ is the diffusion matrix, which is defined via $D_{ij}\delta(t-t^{\prime})=\langle n_{i}(t)n_{j}(t^{\prime})+n_{j}(t^{\prime})n_{i}(t)\rangle/2$. When the CM is obtained, we adopt the logarithmic negativity $E_{N}$ to calculate the degree of any bipartite entanglement of the system, which is defined as [88]
@@
3 Nonreciprocal entanglements in the ideal case